OpenStudy (anonymous):

The leg of a right triangle is 3 units and the hypotenuse is 7 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (wolframwizard):

Use the pythagorean theorem A^2+B^2=C^2. A and B are legs and C is the hypotenuse so plug the numbers given into the equation. 3^2+B^2=7^2. This simplifies into 9+B^2=49. Subtract 9 from both sides of the equation and get B^2=40. This is B^2 and we want to find B so take the square root. B=√40 which is about 6.325 units. If you want the exact value, it is 2√10.